What is the wavelength of the electromagnetic wave
Answer #1


frequency is f = 1/(2*pi*sqrt(L*C)) =1/(2*3.142*sqrt(0.32*10^-6*20.6*10^-12)) = 6.19*10^7 Hz


wave length is lamda = c/f = (3*10^8)/(6.19*10^7) =4.84 m